# LEAF: Unveiling Two Sides of the Same Coin in Semi-supervised Facial Expression Recognition

### v1

@@ Abstract.
Semi-supervised learning has emerged as a promising approach to tackle the challenge of label scarcity in facial expression recognition (FER) task. However, current state-of-the-art methods primarily focus on one side of the coin, i.e., generating high-quality pseudo-labels, while overlooking the other side: enhancing expression-relevant representations. In this paper, we unveil both sides of the coin by proposing a unified framework termed hierarchicaL dEcoupling And Fusing (LEAF) to coordinate expression-relevant representations and pseudo-labels for semi-supervised FER. LEAF introduces a hierarchical expression-aware aggregation strategy that operates at three levels: semantic, instance, and category. (1) At the semantic and instance levels, LEAF decouples representations into expression-agnostic and expression-relevant components, and adaptively fuses them using learnable gating weights. (2) At the category level, LEAF assigns ambiguous pseudo-labels by decoupling predictions into positive and negative parts, and employs a consistency loss to ensure agreement between two augmented views of the same image. Extensive experiments on benchmark datasets demonstrate that by unveiling and harmonizing both sides of the coin, LEAF outperforms state-of-the-art semi-supervised FER methods, effectively leveraging both labeled and unlabeled data. Moreover, the proposed expression-aware aggregation strategy can be seamlessly integrated into existing semi-supervised frameworks, leading to significant performance gains.
Facial expression recognition; semi-supervised learning; Pseudo-labeling; representation learning; consistency regularization
Equal contribution,corresponding author
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Computing methodologies Image representations
1. Introduction
@@
Facial expressions are a critical component of human communication, serving as a primary means of conveying emotions. With the increasing influence of artificial intelligence-generated content, facial expression recognition (FER) has gained significant attention in recent years, finding applications in various domains such as human-machine interaction (Shibata et al., 1997; Sun et al., 2019; Erol et al., 2019) and the development of digital humans (Volonte et al., 2021; Loveys et al., 2021). However, accurately detecting and interpreting facial expressions, particularly in uncontrolled settings, presents substantial challenges for emotion recognition systems.
@@
Figure 1. Our LEAF consistently outperforms state-of-the-art semi-supervised FER approaches across different settings.
@@
The main hurdle in FER stems from the difficulty in obtaining a large volume of labeled facial expression data. This challenge is compounded by high inter-class similarities and the potential for errors, even among well-trained annotators. Consequently, FER under label scarcity has emerged as a practical yet under-explored problem. While most existing FER approaches (Li et al., 2017, 2021a; She et al., 2021; Xue et al., 2021) exhibit a data-hungry nature, heavily relying on extensive labeled data, there is a pressing need for a semi-supervised FER approach (Grandvalet and Bengio, 2004; Rosenberg et al., 2005; Hadsell et al., 2006; Song et al., 2017) that can effectively utilize a small amount of labeled data in conjunction with a large amount of unlabeled data to recognize facial expressions accurately.
@@
Recent progress in semi-supervised FER methods (Li et al., 2022; Florea et al., 2020) and semi-supervised image classification approaches (Sohn et al., 2020; Zhang et al., 2021; Wang et al., 2022; Chen et al., 2023) has primarily focused on one side of the coin: enhancing the quality or quantity of pseudo-labels. These methods aim to generate accurate and diverse pseudo-labels for unlabeled data, which can then be used to train the model in a supervised manner. However, they often overlook the other side of the coin: the potential improvement in representation quality stemming from the inherent minor inter-class differences in FER. Learning expression-relevant representations is crucial, as the differences between various facial expressions can be subtle. The model must be capable of capturing these nuances to accurately classify emotions. By focusing solely on pseudo-labels, existing methods (Song et al., 2017; Hadsell et al., 2006) may fail to learn discriminative representations that can effectively distinguish between different expressions.
@@
In this paper, we propose a novel data-efficient framework called hierarchicaL dEcoupling And Fusing (LEAF) to address both sides of the same coin in semi-supervised FER. LEAF aims to improve the quality of representations as well as the quality and quantity of pseudo-labels from a hierarchical perspective. The core of LEAF lies in gradually teaching the network to distinguish facial expression representations and pseudo-labels into expression-agnostic and expression-relevant parts through decoupling strategies at different levels. The decoupled representations and pseudo-labels are then fused by automatically assigning different weights to them, enabling the model to focus on the parts more relevant to facial expressions, thereby achieving better recognition performance.
@@
Specifically, LEAF introduces three levels of dEcoupling And Fusing (EAF) strategy. At the semantic and instance levels, the EAF strategy draws inspiration from the Mixture-of-Experts (MoE) technique (Eigen et al., 2013; Shazeer et al., 2017; Fedus et al., 2022), allowing individual experts to learn and handle their respective specialized representations (i.e., expression-agnostic and expression-relevant parts). These representations are then automatically weighted and fused through a learnable gating network. At the category level, inspired by existing metric learning works (Liu et al., 2017; Sun et al., 2020; Wang et al., 2018a, b; Qiao et al., 2023), an ambiguous consistency loss is designed to minimize the distance between the prediction distributions obtained in two forward processes. Unlike traditional deterministic consistency losses, LEAF assigns several candidate pseudo-labels to each unlabeled sample, ambiguously labeling them as positive (expression-relevant) or negative (expression-agnostic). A margin is used to control the distance between the positive and negative categories, enhancing the consistency between the distributions.
@@
  • We propose LEAF, a novel semi-supervised FER framework that automatically distinguishes between expression-agnostic and expression-relevant representations and pseudo-labels, and assigns them different weights in a hierarchical decoupling and fusing manner, effectively addressing both sides of the same coin in semi-supervised FER.
@@
2. Related Work
2.1. Facial Expression Recognition
@@
There have been numerous approaches proposed for FER, which can be classified into two primary research categories: methods based on handcrafted features (Hu et al., 2008; Luo et al., 2013; Pietikäinen et al., 2011) and those based on deep learning techniques (Li et al., 2017, 2021a; She et al., 2021; Xue et al., 2021). In traditional research, the emphasis is on extracting texture information from datasets obtained in controlled laboratory settings, such as CK+ (Lucey et al., 2010) and Oulu-CASIA (Zhao et al., 2011). With the advent of large-scale unconstrained FER datasets (Barsoum et al., 2016; Li et al., 2017; Mollahosseini et al., 2017), deep facial expression recognition (DFER) algorithms have emerged, aiming to create effective neural networks or loss functions that can deliver superior performance. For instance, Zhang et al. (Zhang et al., 2024) propose re-balanced attention maps and re-balanced smooth labels to mine extra knowledge from both major and minor classes for imbalanced FER. Wu et al. (Wu and Cui, 2023) leverage facial landmarks to mitigate the impact of label noise in FER.
@@
Most of the existing FER methods are data-hungry. Although several semi-supervised FER approaches (Li et al., 2022; Florea et al., 2020) have been proposed to explore recognizing facial expressions under label scarcity and make some progress, they only concentrate on enhancing the quality or quantity of pseudo-labels, while ignoring that the representations of facial expressions can also be improved.
Figure 2. An overview of LEAF. The weak and strong augmented views of facial expressions are first mapped into the embedding space through a shared encoder. Then we conduct the semantic-level EAF and the instance-level EAF before and after the classifier to reorganize weights for expression-relevant and expression-agnostic representations, respectively. After getting the predictions, we adopt the category-level EAF to generate ambiguous pseudo-labels for consistency regularization.
2.2. Semi-Supervised Learning
@@
In recent years, there has been notable progress in applying semi-supervised learning methods to tackle challenging problems (Sohn et al., 2020; Zhang et al., 2021). These methods utilize various techniques such as consistency regularization (Sajjadi et al., 2016; Xie et al., 2020), entropy minimization (Grandvalet and Bengio, 2004; Lee et al., 2013), and traditional regularization (Berthelot et al., 2019b) to make effective use of unlabeled data. Among them, pseudo-labeling has emerged as a pioneering semi-supervised learning technique for obtaining hard labels from model predictions based. Notably, threshold-based pseudo-labeling approaches have been employed to select unlabeled samples with high-confidence predictions. For instance, some methods (Sohn et al., 2020; Xie et al., 2020) employ a fixed threshold to obtain pseudo-labels and incorporate both weak and strong augmentations to enforce consistency regularization. Other methods (Zhang et al., 2021; Xu et al., 2021) explore dynamic threshold strategies to adaptively determine which samples to assign pseudo-labels.
@@
3. Problem Definition
@@
(1) $\mathcal{L}=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\mathcal{L}^{s}(F(x_{i}^{l};\theta),y_{i})+\frac{\lambda}{N_{u}}\sum_{i=1}^{N_{u}}\mathcal{L}^{u}(F(x_{i}^{u};\theta),\bar{y_{i}}),$
@@
The state-of-the-art SSL approaches for FER (Li et al., 2022; Sohn et al., 2020; Zhang et al., 2021; Chen et al., 2023; Wang et al., 2022) typically formulate the supervised loss as the cross-entropy loss between the model prediction $p_{i}$ and the label $y_{i}$, and the unsupervised loss as the consistency loss between the model prediction $p_{i}$ and the pseudo-label $\bar{y_{i}}$. However, these methods suffer from two main limitations (Sohn et al., 2020; Zhang et al., 2021; Li et al., 2022; Florea et al., 2020): (1) Equal treatment of representations: Existing methods fail to consider the varying discriminative power of different facial expression representations, treating all representations equally. This approach overlooks the fact that some representations may provide more valuable information for accurate recognition than others. (2) Inappropriate pseudo-labeling: The assignment of hard pseudo-labels to unlabeled samples can be problematic, especially when labeled data is scarce. The limited availability of labeled data hinders the model’s ability to make accurate predictions, leading to potentially unreliable pseudo-labels.
4. LEAF Framework
@@
4.1. Semantic-level Decoupling and Fusing
The semantic-level strategy, depicted in Fig. 2 (a), assumes that encoded deep features capture rich geometric or texture information. However, we argue that this information redundancy may hinder fine-grained expression recognition, as not all details contribute equally to the performance of final predictions. Inspired by the powerful Mixture-of-Experts (MoE) technique (Eigen et al., 2013; Shazeer et al., 2017; Fedus et al., 2022), we propose decoupling these features, allowing the network to autonomously determine the usefulness of each feature for expression recognition. As a result, the network learns to assign greater weights to more impactful features.
@@ 4.1. Semantic-level Decoupling and Fusing
(3) $G(f_{i})=\delta(TopK(\sigma(w^{r}\cdot f_{i}))),$
@@
Figure 3. The detailed structure of linear expert, bottleneck expert, and residual expert.
4.2. Instance-level Decoupling and Fusing
@@
(5) $G(p_{i})=\delta(TopK(\sigma(w^{i}\cdot p_{i}))).$
@@
4.3. Category-level Decoupling and Fusing
Several approaches (Sohn et al., 2020; Xie et al., 2020; Zhang et al., 2021; Xu et al., 2021) assign hard pseudo-labels to unlabeled data, aiming to convert unsupervised learning scenarios into supervised ones. However, most of these methods employ a threshold to filter out low-confidence pseudo-labels, utilizing only a portion of unlabeled samples while discarding the rest. Moreover, considering the subtle inter-class differences in FER, fixed pseudo-labels may encounter challenges such as incorrect assignments and discarding useful information from highly similar negative classes.
@@ 4.3. Category-level Decoupling and Fusing
(6) $Y_{sa}^{p}$ =$=\{y_{sa}^{1},y_{sa}^{2},...,y_{sa}^{m}\},$
(7) $Y_{sa}^{n}$ =$=\{y_{sa}^{m+1},y_{sa}^{m+2},...,y_{sa}^{k}\},$
where $m$ denotes the number of positive labels, and the remaining ones are considered as negative labels. We use the cumulative probability of the sorted prediction to determine the value of $m$:
(8) $\sum_{j=1}^{m}\psi(y_{sa}^{j})\geq T,$
where $\psi$ denotes the sort process, and $T$ is a threshold set to $0.9$ empirically. Inspired by metric learning works (Liu et al., 2017; Sun et al., 2020; Wang et al., 2018a, b; Qiao et al., 2023), we force the prediction of weak augmented sample to be consistent with the above partition of the strong augmented sample with a learning objective as:
(9) $min(Y_{wa}^{p})-max(Y_{wa}^{n})>\epsilon\geq 0,$
@@
(11) $min(Y_{wa}^{p})\approx-log(\sum_{i=1}^{m}e^{-y^{i}}),$
(12) $max(Y_{wa}^{n})\approx log(\sum_{j=m+1}^{k}e^{y^{j}}),$
(13) $\eta(Y)=max(Y,0)\approx log(1+e^{Y}).$

### v3

@@ Abstract
Semi-supervised learning has emerged as a promising approach to tackle the challenge of label scarcity in facial expression recognition (FER) task. However, current state-of-the-art methods primarily focus on one side of the coin, i.e., generating high-quality pseudo-labels, while overlooking the other side: enhancing expression-relevant representations. In this paper, we unveil both sides of the coin by proposing a unified framework termed hierarchicaL dEcoupling And Fusing (LEAF) to coordinate expression-relevant representations and pseudo-labels for semi-supervised FER. LEAF introduces a hierarchical expression-aware aggregation strategy that operates at three levels: semantic, instance, and category. (1) At the semantic and instance levels, LEAF decouples representations into expression-agnostic and expression-relevant components, and adaptively fuses them using learnable gating weights. (2) At the category level, LEAF assigns ambiguous pseudo-labels by decoupling predictions into positive and negative parts, and employs a consistency loss to ensure agreement between two augmented views of the same image. Extensive experiments on benchmark datasets demonstrate that by unveiling and harmonizing both sides of the coin, LEAF outperforms state-of-the-art semi-supervised FER methods, effectively leveraging both labeled and unlabeled data. Moreover, the proposed expression-aware aggregation strategy can be seamlessly integrated into existing semi-supervised frameworks, leading to significant performance gains. Our code is available at https://github.com/zfkarl/LEAF.
keywords:
[…]
1 Introduction
@@
Facial expressions are a critical component of human communication, serving as a primary means of conveying emotions. With the increasing influence of artificial intelligence-generated content, facial expression recognition (FER) has gained significant attention in recent years, finding applications in various domains such as human-machine interaction Shibata et al. (1997); Sun et al. (2019); Erol et al. (2019) and the development of digital humans Volonte et al. (2021); Loveys et al. (2021). However, accurately detecting and interpreting facial expressions, particularly in supervised settings, presents substantial challenges for emotion recognition systems.
@@
Figure 1: Our LEAF consistently outperforms state-of-the-art semi-supervised FER approaches across different settings.
@@
The main obstacle in FER stems from the difficulty in obtaining a large volume of labeled facial expression data. This challenge is compounded by high inter-class similarities and the potential for errors, even among well-trained annotators. Consequently, FER under label scarcity has emerged as a practical yet under-explored problem. While most existing FER approaches Li et al. (2017, 2021); She et al. (2021); Xue et al. (2021) exhibit a data-hungry nature, heavily relying on extensive labeled data, there is a pressing need for a semi-supervised FER approach Grandvalet and Bengio (2004); Rosenberg et al. (2005); Hadsell et al. (2006); Song et al. (2017) that can effectively utilize a small amount of labeled data in conjunction with a large amount of unlabeled data to recognize facial expressions accurately.
@@
Recent progress in semi-supervised FER methods Li et al. (2022); Florea et al. (2020) and semi-supervised image classification approaches Sohn et al. (2020); Zhang et al. (2021); Wang et al. (2022); Chen et al. (2023) has primarily focused on one side of the coin: enhancing the quality or quantity of pseudo-labels. These methods aim to generate accurate and diverse pseudo-labels for unlabeled data, which can then be used to train the model in a supervised manner. However, they often overlook the other side of the coin: the potential improvement in representation quality stemming from the inherent minor inter-class differences in FER. Learning expression-relevant representations is crucial, as the differences between various facial expressions can be subtle. The model must be capable of capturing these nuances to accurately classify emotions. By focusing solely on pseudo-labels, existing methods Song et al. (2017); Hadsell et al. (2006) may fail to learn discriminative representations that can effectively distinguish between different expressions.
@@
In this paper, we propose a new data-efficient framework called hierarchicaL dEcoupling And Fusing (LEAF) to address both sides of the same coin in semi-supervised FER. LEAF aims to improve the quality of representations as well as the quality and quantity of pseudo-labels from a hierarchical perspective. The core of LEAF lies in gradually teaching the network to distinguish facial expression representations and pseudo-labels into expression-agnostic and expression-relevant parts through decoupling strategies at different levels. The decoupled representations and pseudo-labels are then fused by automatically assigning different weights to them, enabling the model to focus on the parts more relevant to facial expressions, thereby achieving better recognition performance.
@@
Specifically, LEAF introduces three levels of dEcoupling And Fusing (EAF) strategy. At the semantic and instance levels, the EAF strategy draws inspiration from the Mixture-of-Experts (MoE) technique Eigen et al. (2013); Shazeer et al. (2017); Fedus et al. (2022), allowing individual experts to learn and handle their respective specialized representations (i.e., expression-agnostic and expression-relevant parts). These representations are then automatically weighted and fused through a learnable gating network. At the category level, inspired by existing metric learning works Liu et al. (2017); Sun et al. (2020); Wang et al. (2018a, b); Qiao et al. (2023), an ambiguous consistency loss is designed to minimize the distance between the prediction distributions obtained in two forward processes. Unlike traditional deterministic consistency losses, LEAF assigns several candidate pseudo-labels to each unlabeled sample, ambiguously labeling them as positive (expression-relevant) or negative (expression-agnostic). A margin is used to control the distance between the positive and negative categories, enhancing the consistency between the distributions.
@@
  • We propose a semi-supervised FER framework LEAF that automatically distinguishes between expression-agnostic and expression-relevant representations and pseudo-labels, and assigns them different weights in a hierarchical decoupling and fusing manner, effectively addressing both sides of the same coin in semi-supervised FER.
@@
2 Related Work
2.1 Facial Expression Recognition
@@
There have been numerous approaches proposed for FER, which can be classified into two primary research categories: methods based on handcrafted features Hu et al. (2008); Luo et al. (2013); Pietikäinen et al. (2011) and those based on deep learning techniques Li et al. (2017, 2021); She et al. (2021); Xue et al. (2021). In traditional research, the emphasis is on extracting texture information from datasets obtained in controlled laboratory settings, such as CK+ Lucey et al. (2010) and Oulu-CASIA Zhao et al. (2011). With the advent of large-scale unconstrained FER datasets Barsoum et al. (2016); Li et al. (2017); Mollahosseini et al. (2017), deep facial expression recognition (DFER) algorithms have emerged, aiming to create effective neural networks or loss functions that can deliver superior performance. For instance, Zhang et al. Zhang et al. (2024) propose re-balanced attention maps and re-balanced smooth labels to mine extra knowledge from both major and minor classes for imbalanced FER. Wu et al. Wu and Cui (2023) leverage facial landmarks to mitigate the impact of label noise in FER. Chen et al. Chen et al. (2024) propose to transfer knowledge from static images to unlabeled frames of dynamic videos. Moreover, recent progress in FER lies in human prior-based network Xie et al. (2020); Wang et al. (2021); Li et al. (2021); Gu et al. (2022); Li et al. (2023); Cai et al. (2024), self-supervised learning techniques Liu et al. (2022); Zhang et al. (2024), and cross-modal prompts Zhou et al. (2024).
@@
However, most of the existing FER methods are data-hungry. Although several semi-supervised FER approaches Li et al. (2022); Florea et al. (2020) have been proposed to explore recognizing facial expressions under label scarcity and make some progress, they only concentrate on enhancing the quality or quantity of pseudo-labels, while ignoring that the representations of facial expressions can also be improved.
2.2 Semi-Supervised Learning
@@
In recent years, there has been notable progress in applying semi-supervised learning methods to tackle challenging problems Sohn et al. (2020); Zhang et al. (2021). These methods utilize various techniques such as consistency regularization Sajjadi et al. (2016); Xie et al. (2020), entropy minimization Grandvalet and Bengio (2004); Lee et al. (2013), and traditional regularization Berthelot et al. (2019) to make effective use of unlabeled data. Among them, pseudo-labeling has emerged as a pioneering semi-supervised learning technique for obtaining hard labels from model predictions. Notably, threshold-based pseudo-labeling approaches have been employed to select unlabeled samples with high-confidence predictions. For instance, some methods Sohn et al. (2020); Xie et al. (2020) employ a fixed threshold to obtain pseudo-labels and incorporate both weak and strong augmentations to enforce consistency regularization. Other methods Zhang et al. (2021); Xu et al. (2021) explore dynamic threshold strategies to adaptively determine which samples to assign pseudo-labels.
@@
Figure 2: An overview of LEAF. The weak and strong augmented views of facial expressions are first mapped into the embedding space through a shared encoder. Then we conduct the semantic-level EAF and the instance-level EAF before and after the classifier to reorganize weights for expression-relevant and expression-agnostic representations, respectively. After getting the predictions, we adopt the category-level EAF to generate ambiguous pseudo-labels for consistency regularization.
3 Problem Definition
@@
$\mathcal{L}=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\mathcal{L}^{s}(F(x_{i}^{l};\theta),y_{i})+\frac{\lambda}{N_{u}}\sum_{i=1}^{N_{u}}\mathcal{L}^{u}(F(x_{i}^{u};\theta),\bar{y_{i}}),$ (1)
@@
The state-of-the-art SSL approaches for FER Li et al. (2022); Sohn et al. (2020); Zhang et al. (2021); Chen et al. (2023); Wang et al. (2022) typically formulate the supervised loss as the cross-entropy loss between the model prediction $p_{i}$ and the label $y_{i}$, and the unsupervised loss as the consistency loss between the model prediction $p_{i}$ and the pseudo-label $\bar{y_{i}}$. However, these methods suffer from two main limitations Sohn et al. (2020); Zhang et al. (2021); Li et al. (2022); Florea et al. (2020): (1) Equal treatment of representations: Existing methods fail to consider the varying discriminative power of different facial expression representations, treating all representations equally. This approach overlooks the fact that some representations may provide more valuable information for accurate recognition than others. (2) Inappropriate pseudo-labeling: The assignment of hard pseudo-labels to unlabeled samples can be problematic, especially when labeled data is scarce. The limited availability of labeled data hinders the model’s ability to make accurate predictions, leading to potentially unreliable pseudo-labels.
4 LEAF Framework
@@
4.1 Semantic-level Decoupling and Fusing
The semantic-level strategy, depicted in Fig. 2 (a), assumes that encoded deep features capture rich geometric or texture information. However, we argue that this information redundancy may hinder fine-grained expression recognition, as not all details contribute equally to the performance of final predictions. Inspired by the powerful Mixture-of-Experts (MoE) technique Eigen et al. (2013); Shazeer et al. (2017); Fedus et al. (2022), we propose decoupling these features, allowing the network to autonomously determine the usefulness of each feature for expression recognition. As a result, the network learns to assign greater weights to more impactful features.
@@ 4.1 Semantic-level Decoupling and Fusing
$G(f_{i})=\delta(TopK(\sigma(w^{r}\cdot f_{i}))),$ (3)
@@
Figure 3: The detailed structure of linear expert, bottleneck expert, and residual expert.
4.2 Instance-level Decoupling and Fusing
@@
$G(p_{i})=\delta(TopK(\sigma(w^{i}\cdot p_{i}))).$ (5)
@@
4.3 Category-level Decoupling and Fusing
Several approaches Sohn et al. (2020); Xie et al. (2020); Zhang et al. (2021); Xu et al. (2021) assign hard pseudo-labels to unlabeled data, aiming to convert unsupervised learning scenarios into supervised ones. However, most of these methods employ a threshold to filter out low-confidence pseudo-labels, utilizing only a portion of unlabeled samples while discarding the rest. Moreover, considering the subtle inter-class differences in FER, fixed pseudo-labels may encounter challenges such as incorrect assignments and discarding useful information from highly similar negative classes.
@@ 4.3 Category-level Decoupling and Fusing
$Y_{sa}^{p}$ =$=\{y_{sa}^{1},y_{sa}^{2},...,y_{sa}^{m}\},$ (6)
$Y_{sa}^{n}$ =$=\{y_{sa}^{m+1},y_{sa}^{m+2},...,y_{sa}^{k}\},$ (7)
where $m$ denotes the number of positive emotion classes, and the remaining ones are considered as negative classes. Note that the positive and negative emotion classes here are not pre-defined knowledge in affective computing. Instead, we use the cumulative probability of the sorted prediction to determine the value of $m$:
$\sum_{j=1}^{m}\psi(y_{sa}^{j})\geq T,$ (8)
where $\psi$ denotes the sort process, and $T$ is a threshold set to $0.9$ empirically. Inspired by metric learning works Liu et al. (2017); Sun et al. (2020); Wang et al. (2018a, b); Qiao et al. (2023), we force the prediction of weak augmented sample to be consistent with the above partition of the strong augmented sample with a learning objective as:
$min(Y_{wa}^{p})-max(Y_{wa}^{n})>\epsilon\geq 0,$ (9)
@@
$min(Y_{wa}^{p})\approx-log(\sum_{i=1}^{m}e^{-y^{i}}),$ (11)
$max(Y_{wa}^{n})\approx log(\sum_{j=m+1}^{k}e^{y^{j}}),$ (12)
$\eta(Y)=max(Y,0)\approx log(1+e^{Y}).$ (13)